# AgilePilot: DRL-Based Drone Agent for Real-Time Motion Planning in Dynamic Environments by Leveraging Object Detection

### v1

@@ Abstract
Autonomous drone navigation in dynamic environments remains a critical challenge, especially when dealing with unpredictable scenarios including fast-moving objects with rapidly changing goal positions. While traditional planners and classical optimisation methods have been extensively used to address this dynamic problem, they often face real-time, unpredictable changes that ultimately leads to sub-optimal performance in terms of adaptiveness and real-time decision making. In this work, we propose a novel motion planner, AgilePilot, based on Deep Reinforcement Learning (DRL) that is trained in dynamic conditions, coupled with real-time Computer Vision (CV) for object detections during flight. The training-to-deployment framework bridges the Sim2Real gap, leveraging sophisticated reward structures that promotes both safety and agility depending upon environment conditions. The system can rapidly adapt to changing environments, while achieving a maximum speed of m/s in real-world scenarios. In comparison, our approach outperforms classical algorithms such as Artificial Potential Field (APF) based motion planner by times, both in performance and tracking accuracy of dynamic targets by using velocity predictions while exhibiting success rate in 75 conducted experiments. This work highlights the effectiveness of DRL in tackling real-time dynamic navigation challenges, offering intelligent safety and agility.
@@ I Introduction
In recent years, autonomous drone navigation has gained significant interest, especially in the context of dynamic environments. Conventional motion planning methods, such as nonlinear Model Predictive Control (NMPC) and graph-based search algorithms, have demonstrated efficiency in structured and controlled environments but lack their capability to adapt to rapidly changing conditions in real-time. Therefore, DRL has emerged as a viable alternative, offering fast decision-making and adaptability that improve autonomous navigation. In recent work, Song et al. [1] has explored near-optimal trajectory generation in near time using DRL methods, enabling adaptation to dynamic track configurations while achieving remarkable speeds of up to 60 km/h. Another study [2] by Kaufmann et al. utilises Model Predictive Control (MPC) to navigate quickly through the track, where a convolutional network and an Extended Kalman Filter (EKF) predicts the poses of the closest gates along with uncertainty. Moreover, another breakthrough study [3] presents a SWIFT system that competes with human drone pilots in drone racing by using the Proximal Policy Optimisation (PPO) algorithm to train a model that generates low-level control commands for the quadrotor. Despite these studies having developed agile systems for drones, they lack the ability to effectively avoid obstacles in real time and adapt to a highly dynamically changing environment. In addition, there are also limitations to their ability to balance safety and agility when the environment becomes highly unpredictable.
This research introduces a novel motion planner, AgilePilot, which leverages DRL to predict drone velocities for navigating dynamic environments with moving objects. The velocity prediction is situation-aware reaching up to 3.0 m/s in simpler terrain while slowing down in complex scenarios involving obstacles and moving objects. The proposed methodology involves a unique framework to train models in a simulation environment customized to achieve the level of accuracy required for Sim2Real transfer. The training is performed using Actor Critic neural network architecture and uses a unique reward structure with effective randomisation during episodes to make the application of trained models generalised and easy to implement. To track movable objects, the system applies YOLO-based object detection combined with state estimation using IPPE PnP and an EKF for robust pose estimation.
The methods used in this research revolutionise the motion planning of drones in dynamic environments by effectively adapting the velocity vector according to the complexity of terrain while taking account uncertainty in the objects movements. The broad applicability of technology is ensured by the pipeline used for training and testing of trained models in unknown environments.
@@ II Related Works
Due to the overwhelming complexities of drone navigation in dynamic environments, many studies have been trying to solve the general applicability of algorithms either by using machine learning or classical methods. For example, in study [4], the policy was learnt with high uncertainty and tested in a noisy real environment. The results showed effective drone navigation in reality, but the method is applicable in 2D space only and not tested in dynamic conditions.
@@
In study [5], DRL is utilised to avoid moving and stationary obstacles consisting of images and several scalars with Joint Neural Network (JNN). Liu et al., in research [6] use a motion planning framework that integrates visibility path searching to generate collision-free paths and RL to generate low-level motion commands. Moreover, CPU-based trained DRL for UAV to UAV tracking is introduced in [7]. The model was trained using PPO and showed considerable results. Another PPO-based learnt RL agent in study [8] was used in a racing environment as a path planner and utilized conventional Proportional–Integral–Derivative (PID) controller for controlling the motion. Moreover, Song et al. research [9] generates smooth collision-free trajectories while taking care of an unseen environment by incorporating vision with DRL. While these mentioned studies show promising capabilities for motion generation and obstacle avoidance, they have predominantly been tested in simulation environments only with less challenging scenarios.
@@
Furthermore, in study [10] highly efficient path planner was devised for cluttered environments incorporating some extent of dynamical feasibility by adjusting time allocation based upon spatial-temporal joint optimisation. Moreover, study [11] introduced a search-based method by using a multi-stage training approach and testing the agent on real air-ground unmanned system. In [12], DRL was trained for random obstacles in dense urban environment by utilising Double Deep-Q network (DDQP) while enhancing the training stability of algorithms. In addition, study [13] improves safe navigation in obstacle avoidance by leveraging map parameterisation and low-cost planning whereas the results are validated in real-life experiments. Peter et. al in study [14] trained models using Twin-Delayed Deep Deterministic (TD3) policy gradient algorithm and real life experiments showed remarkable landing performance of drones for dynamically moving platforms in the presence of downwash and wind turbulence. The extension of these works using PPO has also been successfully implemented for multi-agents [15]. However, these approaches focus purely on path planning without considering obstacles and have only been tested on tiny-sized drones, therefore, overall restricting the high velocity potential of medium-sized drones. Lastly, another study [16] introduced OmniRace, a control interface based on drone velocity manipulation for drones of all sizes. However, its motion planning depends entirely on human input, which limits both navigation accuracy and autonomy.
@@
In order to address the mentioned gaps, this research presents a novel DRL-based motion planning framework that enables velocity-based trajectory generation for any size drone due to its nature of being a model-free learning agent. Unlike previous studies that rely on static environment, our approach is designed for dynamic scenarios, ensuring robust adaptabiltiy to rapidly changing surroundings. The training environment used is Gym PyBullet [17] and carefully customised and tuned to incorporate real-world complexities for a seamless transition from simulation to reality. Furthermore, this framework accounts for dynamically evolving obstacles, allowing the drone to make intelligent decisions in real time while balancing agility and safety as required.
@@ III Methodology
This section explains the methods used for developing AgilePilot pipeline. Gym PyBullet is used for simulation and physics in training with PPO and models are deployed in simulated and real-life environments. The overall pipeline of our system is shown in Fig. 2.
@@ III-A Simulation Environment
The simulation setup is developed using Gym environment and PyBullet physics engine that allows realistic aerodynamics and accurate quadcopter mathematical modelling to emulate drones. We have integrated a custom midsized drone with a custom PID controller in simulation. The inertial and controller parameters are tuned to the level of realistic response in real environments. Moreover, custom objects such as gates and obstacles are also built, as shown in the simulation Fig. 3.
In each step, the object’s movement is governed by a random velocity drawn from a uniform probability distribution.
Where represents the change in position in the and directions, and denotes a uniform distribution over the range to .
Lastly, the PID controller allows drone to reach speed within range of m/s in and in direction whereas rotation angles can range from to radians, by mapping desired commands to motor RPMs and subsequently to desired forces and torques by using accurate motor dynamics model.
@@ III-B1 Neural Network Architecture
The output layer of actor represents the action space, which consists of the following four values:
@@ III-B2 Reward Structure
The agent receives a reward based on its distance to the goal. The reward increases substantially when the drone is in close proximity to the goal which allows stable target location tracking in 3D space.
where is the Euclidean distance between the drone’s position and the goal and is the constant to avoid extremely high reward values.
Moreover, the agent is penalised for getting closer to the obstacle where the obstacles are modelled as infinite-height objects. The penalty is based on the relative distance to the obstacle and its size, using an exponential function:
where equals , is the penalty scaling constant, and is the safety region around the obstacle.
@@
where is the Euclidean norm of the velocity vector and is the velocity penalty constant.
@@ III-B3 Training
The obstacle position is randomized based on two components: longitudinal and lateral offsets.
The longitudinal offset is calculated along the line connecting the drone and the target, while the lateral offset is perpendicular to this line. Let be the vector from the drone to the target:
@@
The obstacle’s position is then determined by combining the longitudinal and lateral offsets with drone’s initial position in the episode. The -coordinate of the obstacle is randomised within a range, typically around .
@@
The feature extractor (policy network) processes observations and provides the necessary features for action prediction. For PPO, the policy architecture is based on an ActorCriticPolicy which integrates both the actor (policy) and critic (value function) into a single neural network. The output layer provides a distribution of actions based on the processed input.
The model’s performance is evaluated after training by running it in the environment and computing the mean and standard deviation of the rewards over multiple episodes. Fig. 6 and Fig. 7 shows mean reward and mean episode length, respectively of several trained models with different set of hyperparameters. Model 2 with hyperparmeters listed in Table. I was selected for experiment due to its better convergence with highest reward and minimum episode length indicating that the policy has learned well.
@@ III-C1 Object Detection
The object detection system receives an image from the wide-angle Intel RealSense T265 camera with 30 fps as input and outputs an estimate of the spatial position of the gates and obstacles relative to the camera. The camera captures a frame, which is then processed by the YOLOv8n Pose neural detector. The detector identifies four key points that define the gate plane with size 1.5 x 1.5 m, as well as three key points for cylindrical obstacle with length of 1.0 m and diameter of 0.1 m. To train the detection model, a custom dataset was collected, consisting of 2,475 annotated instances of racing gates and 3,015 instances of obstacles. 4,471 images were generated from the original data set using specialized augmentation techniques. The final dataset was split into training, and validation sets in a 80:20 ratio. The YOLO model was trained for 500 epochs with a batch size of 32 with image resolution 424x400 px. Data augmentation was applied, including horizontal flipping with a probability of 0.5, random rotation up to 10 degrees, and color adjustments in the HSV space with the following parameters: hue 0.015, saturation 0.2, and value 0.2. The inference time of the model is 40 ms, demonstrating the system’s capability for real-time processing.
@@ III-C2 Position estimation
The Infinitesimal Plane-based Pose (IPPE) method facilitates the estimation of the spatial position and orientation of racing gates and obstacles relative to the drone’s camera. This is accomplished using the 2D coordinates of the object’s corner points, the 3D coordinates of these corners within the object’s local coordinate system and the intrinsic camera parameters. The method allows for the precise determination of the pose and orientation of the object in space. The EKF is employed to filter noisy position measurements of multiple objects, enabling accurate tracking by associating observations based on their spatial proximity. The state vector for gates is represented as and for cylindrical obstacle represent the position and orientation of an object, where: are the position coordinates, and is the yaw angle (orientation). The results of the working detection and pose estimation system are presented in Fig. 8.
@@ IV Comparative Analysis
In this analysis, we compare the performance of our DRL-based motion planner with the well-known Artificial Potential Field (APF) motion planner due to its effectiveness in dynamic environments for UAVs [18] [19]. To perform this comparison, we conducted five simulation cases to assess different dynamic conditions. The general configuration for each case includes two obstacles and one gate that the drone must pass through to reach the target location.
@@ IV-A Simulation Cases
Case 1: Moving gate at 0.3 m/s with obstacle placed to exploit APF local minima.
Case 2: Moving gate only at 0.3 m/s
Case 3: Moving gate at 0.3 m/s with varying target height.
Case 4: Both moving gate and obstacle at 0.3 m/s with varying target height.
Case 5: Gate moving only at 0.6 m/s.
Each case is performed 15 times with distinctive initial drone and object positions to record the number of failures where failure exists whenever the drone is collided with the object in environment.
@@ IV-B Performance Metrics
Success Rate: The success rate is calculated using the formula:
@@
Table II showcases average success rate out of total 75 experiments by DRL agent as compared to APF motion planner where APF completely failed to navigate through local minima situation in case 1 and struggled to complete its task for highly dynamically moving environment in case 4 and case 5.
@@
Furthermore, Table III indicates key performance parameters between the two approaches. DRL agent is able to perform high precision gate and target tracking via velocity predictions i.e., approximately 3.0 times better than APF planner. Moreover, the time required to complete all the cases on average is 5.0 seconds which is approximately 3.0 times faster than APF planning in dynamic environments.
From one of the visual representation of the cases, high precision with high speed navigation is also evident from Fig. 9 where it can be seen that the drone follows through the moving gate rather smoothly and intelligently where it speeds goes beyond 2.5 m/s. On the contrary, APF planner struggles to adapt to the moving gate scenario and maintain its speed below 1.0 m/s in doing so.

### v2

@@ Abstract
Autonomous drone navigation in dynamic environments remains a critical challenge, especially when dealing with unpredictable scenarios including fast-moving objects with rapidly changing goal positions. While traditional planners and classical optimization methods have been extensively used to address this dynamic problem, they often face real-time, unpredictable changes that ultimately lead to suboptimal performance in terms of adaptiveness and real-time decision-making. In this work, we propose a novel motion planner, AgilePilot, based on deep reinforcement learning (DRL) that is trained in dynamic conditions, coupled with real-time computer vision (CV) for object detections during flight. The training-to-deployment framework bridges the Sim2Real gap, leveraging sophisticated reward structures that promote both safety and agility depending upon environmental conditions.
The system can rapidly adapt to changing environments while achieving a maximum allowable speed of m/s in real-world scenarios. In comparison, our approach outperforms classical algorithms such as the Artificial Potential Field (APF)-based motion planner by times, both in performance and tracking accuracy of dynamic targets by using velocity predictions while exhibiting a success rate in conducted experiments. This work highlights the effectiveness of DRL in tackling real-time dynamic navigation challenges, offering intelligent safety and agility.
@@ I Introduction
In recent years, autonomous drone navigation has gained significant interest, especially in the context of dynamic environments. Conventional motion planning methods, such as nonlinear Model Predictive Control (NMPC) and graph-based search algorithms, have demonstrated efficiency in structured and controlled environments but lack the capability to adapt to rapidly changing conditions in real-time. Therefore, DRL has emerged as a viable alternative, offering fast decision-making and adaptability that improve autonomous navigation. In recent work, Song et al. [1] have explored near-optimal trajectory generation in near time using DRL methods, enabling adaptation to dynamic track configurations while achieving remarkable speeds of up to 60 km/h. Another study [2] by Kaufmann et al. utilizes Model Predictive Control (MPC) to navigate quickly through the track, where a convolutional network and an Extended Kalman Filter (EKF) predict the poses of the closest gates along with uncertainty. Moreover, another breakthrough study [3] presents a SWIFT system that competes with human drone pilots in drone racing by using the Proximal Policy Optimization (PPO) algorithm to train a model that generates low-level control commands for the quadrotor. Despite these studies having developed agile systems for drones, they lack the ability to effectively avoid obstacles in real time and adapt to a highly dynamically changing environment. In addition, there are also limitations to their ability to balance safety and agility when the environment becomes highly unpredictable.
This research introduces a novel motion planner, AgilePilot, which leverages DRL to predict drone velocities for navigating dynamic environments with moving objects. The velocity prediction is situation-aware, reaching up to 3.0 m/s in simpler terrain while slowing down in complex scenarios involving obstacles and moving objects. The proposed methodology involves a unique framework to train models in a simulation environment customized to achieve the level of accuracy required for Sim2Real transfer. The training is performed using actor-critic neural network architecture and uses a unique reward structure with effective randomization during episodes to make the application of trained models generalized and easy to implement. To track movable objects, the system applies YOLO-based object detection combined with state estimation using IPPE PnP and an EKF for robust pose estimation.
The methods used in this research revolutionize the motion planning of drones in dynamic environments by effectively adapting the velocity vector according to the complexity of terrain while taking into account uncertainty in the objects movements. The broad applicability of technology is ensured by the pipeline used for training and testing of trained models in unknown environments.
@@ II Related Works
Due to the overwhelming complexities of drone navigation in dynamic environments, many studies have been trying to solve the general applicability of algorithms either by using machine learning or classical methods. For example, in study [4], the policy was learned with high uncertainty and tested in a noisy real environment. The results showed effective drone navigation in reality, but the method is applicable in 2D space only and not tested in dynamic conditions.
@@
In study [5], DRL is utilized to avoid stationary and moving obstacles, with observation consisting of images and several scalars with Joint Neural Network (JNN). Liu et al. [6] use a motion planning framework that integrates visibility path searching to generate collision-free paths and RL to generate low-level motion commands. Moreover, CPU-based trained DRL for UAV-to-UAV tracking is introduced in [7]. The model was trained using PPO and showed considerable results. Another PPO-based learned RL agent in study [8] was used in a racing environment as a path planner and utilized a conventional proportional–integral–derivative (PID) controller for controlling the motion. Moreover, Song et al. [9] proposed a system that generates smooth, collision-free trajectories while taking care of an unseen environment by incorporating vision with DRL. While these mentioned studies show promising capabilities for motion generation and obstacle avoidance, they have predominantly been tested in simulation environments only with less challenging scenarios.
@@
Furthermore, in study [10] highly efficient path planner was devised for cluttered environments incorporating some extent of dynamical feasibility by adjusting time allocation based upon spatial-temporal joint optimization. Moreover, study [11] introduced a search-based method by using a multi-stage training approach and testing the agent on a real air-ground unmanned system. In [12], DRL was trained for random obstacles in a dense urban environment by utilizing a Double Deep-Q network (DDQP) while enhancing the training stability of algorithms. In addition, study [13] improves safe navigation in obstacle avoidance by leveraging map parameterization and low-cost planning, where the results are validated in real-life experiments. Peter et al. in study [14] trained models using the Twin-Delayed Deep Deterministic (TD3) policy gradient algorithm, and real-life experiments showed remarkable landing performance of drones for dynamically moving platforms in the presence of downwash and wind turbulence. The extension of these works using PPO has also been successfully implemented for multi-agents [15]. However, these approaches focus purely on path planning without considering obstacles and have only been tested on tiny-sized drones, therefore, overall restricting the high velocity potential of medium-sized drones. Lastly, another study [16] introduced OmniRace, a control interface based on drone velocity manipulation for drones of all sizes. However, its motion planning depends entirely on human input, which limits both navigation accuracy and autonomy.
@@
In order to address the mentioned gaps, this research presents a novel DRL-based motion planning framework that enables velocity-based trajectory generation for any size drone due to its nature of being a model-free learning agent. Unlike previous studies that rely on static environments, our approach is designed for dynamic scenarios, ensuring robust adaptability to rapidly changing surroundings. The training environment used is Gym PyBullet [17] and is carefully customized and tuned to incorporate real-world complexities for a seamless transition from simulation to reality. Furthermore, this framework accounts for dynamically evolving obstacles, allowing the drone to make intelligent decisions in real time while balancing agility and safety as required.
@@ III Methodology
This section explains the methods used for developing the AgilePilot pipeline. Gym PyBullet is used for simulation and physics in training with PPO, and models are deployed in simulated and real-life environments. The overall pipeline of our system is shown in Fig. 2.
@@ III-A Simulation Environment
The policy is trained using an actor-critic network utilizing the PPO algorithm in a customized simulated environment. The position estimation uses a VICON tracking system to provide the drone’s state, while object detection is performed using a YOLO neural network for gate corner detection and obstacle keypoint extraction. The keypoints are processed using IPPE PnP to estimate local positions, which are then mapped to a 3D pose in the global coordinate system. The simulation setup is developed using the Gym environment and the PyBullet physics engine that allows realistic aerodynamics and accurate quadcopter mathematical modeling to emulate drones. We have integrated a custom midsized drone with a custom PID controller in simulation. The inertial and controller parameters are tuned to the level of realistic response in real environments. Moreover, custom objects such as gates and obstacles are also built, as shown in the simulation in Fig. 3.
In each step, the object’s movement is governed by a random velocity drawn from a uniform probability distribution:
where is the change in position in the and directions, and is the uniform distribution over the range to .
Lastly, the PID controller allows the drone to reach velocities within a range of m/s in and in direction, whereas rotation angles can range from to radians, by mapping desired commands to motor RPMs and subsequently to desired forces and torques by using an accurate motor dynamics model.
@@ III-B1 Neural Network Architecture
The output layer of the actor represents the action space, which consists of the following four values:
@@ III-B2 Reward Structure
The agent receives a reward based on its distance to the goal. The reward increases substantially when the drone is in close proximity to the goal, which allows stable target location tracking in 3D space.
where is the Euclidean distance between the drone’s position and the goal, and is the constant to avoid extremely high reward values.
Moreover, the agent is penalized for getting closer to the obstacle where the obstacles are modelled as infinite-height objects. The penalty is based on the relative distance to the obstacle and its size, using an exponential function:
where equals , is the penalty scaling constant and is the safety region around the obstacle.
@@
where is the Euclidean norm of the velocity vector, and is the velocity penalty constant.
@@ III-B3 Training
The obstacle position is randomized based on two components: longitudinal and lateral offsets. The longitudinal offset is calculated along the line connecting the drone and the target, while the lateral offset is perpendicular to this line. is the vector from the drone to the target, calculated as:
@@
The obstacle’s position is then determined by combining the longitudinal and lateral offsets with the drone’s initial position in the episode. The -coordinate of the obstacle is randomized within a range, typically around .
@@
The feature extractor (policy network) processes observations and provides the necessary features for action prediction. For PPO, the policy architecture is based on anactor-critic, which integrates both the actor (policy) and critic (value function) into a single neural network. The output layer provides a distribution of actions based on the processed input.
The model’s performance is evaluated after training by running it in the environment and computing the mean and standard deviation of the rewards over multiple episodes. Fig. 6 and Fig. 7 show the mean reward and mean episode length, respectively, of several trained models with different sets of hyperparameters. Model 2 with hyperparameters listed in Table. I was selected for the experiment due to its better convergence with the highest reward and minimum episode length, indicating that the policy has learned well.
@@ III-C1 Object Detection
The object detection system receives an image from the wide-angle Intel RealSense T265 camera with 30 fps as input and outputs an estimate of the spatial position of the gates and obstacles relative to the camera. The camera captures a frame, which is then processed by the YOLOv8n Pose neural detector. The detector identifies four key points that define the gate plane with a size of 1.5 x 1.5 m, as well as three key points for the cylindrical obstacle with a length of 1.0 m and a diameter of 0.1 m. To train the detection model, a custom dataset was collected, consisting of 2475 annotated instances of racing gates and 3015 instances of obstacles. 4471 images were generated from the original data set using specialized augmentation techniques. The final dataset was split into training and validation sets in an 80:20 ratio. The YOLO model was trained for 500 epochs with a batch size of 32 with an image resolution of 424x400 px. Data augmentation was applied, including horizontal flipping with a probability of 0.5, random rotation up to 10 degrees, and color adjustments in the HSV space with the following parameters: hue 0.015, saturation 0.2, and value 0.2. The inference time of the model is 40 ms, demonstrating the system’s capability for real-time processing.
@@ III-C2 Position estimation
The Infinitesimal Plane-based Pose (IPPE) method facilitates the estimation of the spatial position and orientation of racing gates and obstacles relative to the drone’s camera. This is accomplished using the 2D coordinates of the object’s corner points, the 3D coordinates of these corners within the object’s local coordinate system and the intrinsic camera parameters. The method allows for the precise determination of the pose and orientation of the object in space. The EKF is employed to filter noisy position measurements of multiple objects, enabling accurate tracking by associating observations based on their spatial proximity. The state vector for gates is represented as and for cylindrical obstacle represent the position and orientation of an object, where: are the position coordinates and is the yaw angle (orientation). The results of the working detection and pose estimation system are presented in Fig. 8.
@@ IV Comparative Analysis
In this analysis, we compare the performance of our DRL-based motion planner with the well-known Artificial Potential Field (APF) motion planner due to its effectiveness in dynamic environments for UAVs [18], [19]. To perform this comparison, we conducted five simulation cases to assess different dynamic conditions. The general configuration for each case includes two obstacles and one gate that the drone must pass through to reach the target location.
@@ IV-A Simulation Cases
Case 1: Gate moving at 0.3 m/s with obstacle placed to exploit APF local minima.
Case 2: Gate moving at 0.3 m/s.
Case 3: Gate moving at 0.3 m/s, varying target height.
Case 4: Gate and obstacle moving at 0.3 m/s, varying target height.
Case 5: Gate moving at 0.6 m/s.
Each case is performed 15 times with distinctive initial drone and object positions to record the number of failures, where failure exists whenever the drone collides with the object in the environment.
@@ IV-B Performance Metrics
Success Rate: The success rate is calculated as:
@@
Table II showcases the average success rate out of a total of 75 experiments by the DRL agent as compared to the APF motion planner, where the APF completely failed to navigate through the local minima situation in Case 1 and struggled to complete its task for the highly dynamically moving environment in Case 4 and Case 5.
@@
Furthermore, Table III indicates key performance parameters between the two approaches: mean error (ME), standard deviation (SD), and time taken (TT) to reach the target point for only successful flights. The DRL agent is able to perform high-precision gate and target tracking via velocity predictions, i.e., approximately 3.0 times better than the APF planner. Moreover, the time required to complete all the cases on average is 5.0 s, which is approximately 3.0 times faster than APF planning in dynamic environments.
From one of the visual representations of the cases, high precision with high-speed navigation is also evident from Fig. 9 where it can be seen that the drone follows through the moving gate rather smoothly and intelligently, where its speed goes beyond 2.5 m/s. On the contrary, the APF planner struggles to adapt to the moving gate scenario and maintain its speed below 1.0 m/s in doing so.